# Too Hot to Handle: Searching for Inflationary Particle Production in Planck Data

### v1

@@ V Application to Planck
Finally, we apply our pipeline to the observed Planck dataset. For this purpose, we use the Planck PR4 HFI temperature maps [20, 43, 44, 45], in combination with the point-source and Galactic masks described above. We search for hotspots using $100$ templates across $10$ logarithmically-spaced values of $\eta_{\star}$ and $10$ linearly-spaced values of $\eta_{\rm HS}$, obtaining a catalog of detection candidates as before. This requires $\sim 250$ CPU-hours in total. To account for any candidates appearing multiple times in the catalog, we merge any pair separated by less than the mean hotspot radius ($\approx\sqrt{4\pi}(\eta_{\star}^{1}+\eta_{\star}^{2})/2\eta_{0})$, [17]). To validate any potential detections, we additionally perform an analysis using the sevem and commander component-separated maps [20]. This proceeds as above, but effectively involves only a single-frequency matched filter rather than MMF, and additionally uses the Planck common component-separation mask [46].
Figure 4: Visualuzation of the hotspot detection candidates from the full Planck PR4 high-frequency data. The dark and light shaded regions show the region masked out by the analysis mask and post-processing mask respectively. Each circle represents a single detection candidate with $\mathrm{SNR}>5$, whose size is proportional to the SNR and whose color indicates the inferred value of $\eta_{\star}$. The dark purple circle in the south-west corresponds to multiple similarly-located candidates; as discussed in the text, this is likely due to an unsubtracted compact source. A histogram of the detection significances is shown in Fig. 5, with the parameters of the highest-SNR detections given in Tab. 1.
In Fig. 4, we show the distribution of $\mathrm{SNR}>5$ hotspot candidates across the sky. Outside the mask, these appear roughly isotropically distributed, with an array of different $\eta_{\star}$ values. This is good: we do not find that all detection candidates cluster in regions of higher Galactic contamination. A more useful diagnostic is the distribution of SNR values, which is shown in Fig. 5. In total, we find $144$ candidates, with SNR up to $8$; it is important to note that our analysis has four free parameters ($\eta_{\star}$, $\eta_{\rm HS}$, and the hotspot center coordinates), thus moderately high SNR values are expected due to chance fluctuations. Furthermore, analysis of the sevem (commander) component-separated map finds only 48 (35) candidates, with a maximum detection significance of $6\sigma$. This indicates that the $\mathrm{SNR}>6$ candidates (which all have $\mathrm{SNR}<5$ in component-separated maps) may be spurious instead of primordial.
@@
Tab. 1 lists the parameters of the six-frequency Planck $\mathrm{SNR}>6$ hotspot candidates, with visualizations of the Planck maps around each point shown in Fig. 7. Each corresponds to a hotspot with temperature perturbation around $30\,\,\mu$K. Notably, all potential hotspots are small, with $\eta_{\star}<30\,\,\mathrm{Mpc}$ in all cases; as shown previously, physical hotspots of this size are the hardest to detect due to the beam and noise properties. Furthermore, six of the ten sources are located in a ring of radius $\sim 1^{\circ}$ and have similar inferred parameters, which suggests a ringing signature around a correlated source. Indeed, cross-matching with the public Planck catalogs (including point sources, Galactic cold clumps, tSZ clusters, and high-redshift sources), we find a non-thermal source (i.e., synchrotron) [47] at the center of the ring, which is clearly seen in Fig. 7. We additionally find a non-thermal source within $10^{\prime}$ of the $\mathrm{SNR}=6.0$ candidate and a $5\sigma$ point source within $13^{\prime}$ of the $\mathrm{SNR}=6.9$ candidate. All remaining sources are excluded by the Planck GAL020 mask (though this represents a harsh foreground cut), with the $\mathrm{SNR}=6.2$ candidate masked in the Planck inpainting map. All-in-all, we find no significant evidence for a hotspot with $\mathrm{SNR}>6$ in the Planck data.
Figure 5: PDF of Planck PR4 detections at $\mathrm{SNR}>5$. We compare results from the full six HFI channels (blue) and those from component-separated maps (orange, green) using two pipelines. These detections appear compatible with noise, given that the $\mathrm{SNR}>6$ detections are not found in the component-separated pipeline.
@@ VI Summary & Discussion
Searching for localized features in the CMB is a task well-known to cluster cosmologists. Taking advantage of this parallel, we have utilized codes developed for thermal Sunyaev-Zel’dovich analysis to search for particle-production signatures in Planck PR4 data using matched-filter methods. It is difficult to overemphasize the simplicity of this approach: one simply replaces a tSZ cluster profile (and tSZ frequency dependence) with the desired primordial template (and blackbody frequency dependence). Extensive verification has demonstrated that this approach works: we can recover the input parameters of inflationary hotspots (both singlets and pairs) injected in realistic simulations, given sufficiently large couplings. From the Planck data, we find no evidence for new physics, with no robust detection of hotspots with $\mathrm{SNR}>6$ (using a high threshold due to the four free parameters in the model). This yields a bound on shift-symmetry-violating couplings between the inflaton and a much heavier field, with $g\lesssim 10$ at time $\eta_{\star}>100\,\,\mathrm{Mpc}$. Assuming $60$ $e$-foldings of inflation between the horizon exit of a mode with physical size $1/H_{0}$ (the present day inverse Hubble parameter) and the end of inflation, $\eta_{\star}=100$ Mpc corresponds to $\approx 56$ $e$-foldings before the end of inflation.
@@
An important feature of the profile-finding analyses of this work is that their sensitivity does not depend on the number of hotspots present; we can detect any hotspot given sufficiently large $g,\eta_{\star}$. This differs significantly from power-spectrum-based analyses, whereupon the modification to $C_{\ell}^{TT}$ scales as $g^{2}N_{\rm spots}$ [e.g., 17]. In the right panel of Fig. 6, we compare the $5\sigma$ bound on $g$ from the two approaches, considering both one and one hundred hotspots. In the former case, the constraints from the localized profile-finding analysis dwarf those from the power spectrum, whilst in the latter case, searching for modifications to $C_{\ell}^{TT}$ may yield improved bounds, particularly at high $\eta_{\star}$. This indicates an important and generic point: rare events are best searched for via local-in-space analyses, whilst common events (giving some spectrum of sources) can be probed best with low-order correlators.
The analysis developed herein can be extended in a number of ways. Firstly, we have considered only temperature anisotropies; the Planck CMB dataset also contains polarization information, which can also be used to probe primordial hotspots since our mechanism produces adiabatic fluctuations that contribute to both $T$ and $E$. This will be particularly relevant for future surveys since large-scale temperature anisotropies are already cosmic-variance-limited. Furthermore, high-resolution data from the Atacama Cosmology Telescope [49, 50], South Pole Telescope [51, 52], Simons Observatory [53], and CMB-S4 [54] will facilitate analysis of smaller hotspots, which are indistinguishable from point sources at Planck resolution. This is shown in the right panel of Fig. 6, which gives the upper bounds on $g$ from an idealized experiment with neither a beam nor experimental noise for $\ell\in[2,3500]$. As shown, one could improve the limit on $g$ for small $\eta_{\star}$ by at least an order of magnitude in the future, though cosmic variance limits improvements at large $\eta_{\star}$. Finally, we note that our pipeline can be simply extended to a wide variety of other primordial features, for example those of [14, 19], allowing new insights into physics at the highest energies.
@@ Acknowledgements.
We thank Taegyun Kim and Neal Weiner for insightful comments and discussions. OHEP is a Junior Fellow of the Simons Society of Fellows and thanks Bandai Namco Entertainment Inc. for inspiring the NYC particle astrophysics and cosmology group meeting, as well as Netflix for motivating the title. JCH acknowledges support from NSF grant AST-2108536, DOE grant DE-SC0011941, NASA grants 21-ATP21-0129 and 22-ADAP22-0145, the Sloan Foundation, and the Simons Foundation. This work utilized numpy [55], matplotlib [56], healpy [57], and HEALPix [58].
@@ Appendix A Visual Inspection of Planck Hotspot Candidates
In Fig. 7, we plot Cartesian projections of the Planck HFI frequency maps centered on the $\mathrm{SNR}>6$ hotspot candidates given in Tab. 1. We also show the SNR map (given by $\hat{g}/\sigma(g)$, as defined in Eq. (6) with optimal $\eta_{\star},\eta_{\rm HS}$), computed for the $14.8^{\circ}\times 14.8^{\circ}$ tile in which the candidate was found. The first column appears consistent with a clump of dust (seen clearly at high frequency). Furthermore, we see a ringing feature for the third, which circles a synchrotron source identified in [47]. For the other sources, visual inspection does not yield definitive results.
Figure 7: SNR maps (top row) and frequency maps (other rows, with frequency indicated in GHz) displaying the $\mathrm{SNR}>6$ hotspot candidates, which are circled in green. Each image is a $5^{\circ}\times 5^{\circ}$ Cartesian projection around the centers given in the title. The SNR maps have colorbar truncated to ±$\pm 5$, whilst the maximal value for each other plot is shown in purple. The bounding rectangles demarcate the flat-sky tiles used in the analysis, and the holes are due to point-source masking.

### v2

@@ V Application to Planck
Finally, we apply our pipeline to the observed Planck dataset. For this purpose, we use the Planck PR4 HFI temperature maps [21, 46, 47, 48], in combination with the point source and Galactic masks described above. We search for hotspots using $100$ templates across $10$ logarithmically-spaced values of $\eta_{\star}$ and $10$ linearly-spaced values of $\eta_{\rm HS}$, obtaining a catalog of detection candidates as before. This requires $\sim 250$ CPU-hours in total. To account for any candidates appearing multiple times in the catalog, we merge any pair separated by less than the mean hotspot radius ($\approx\sqrt{4\pi}(\eta_{\star}^{1}+\eta_{\star}^{2})/2\eta_{0})$, [18]). To validate any potential detections, we additionally perform an analysis using the sevem and commander component-separated maps [21]. This proceeds as above, but effectively involves only a single-frequency matched filter rather than MMF, and additionally uses the Planck common component-separation mask, which excludes any point sources with $\mathrm{SNR}\geq 5$ [49].
Figure 4: Visualuzation of the hotspot detection candidates from the full Planck PR4 high-frequency data. The dark and light shaded regions show the region masked out by the analysis mask and post-processing mask respectively. Each circle represents a single detection candidate with $\mathrm{SNR}\geq 5$, whose size is proportional to the SNR and whose color indicates the inferred value of $\eta_{\star}$. The dark purple circle in the south-west corresponds to multiple similarly-located candidates; as discussed in the text, this is likely due to an unsubtracted compact source. A histogram of the detection significances is shown in Fig. 5, with the parameters of the highest-SNR detections given in Tab. 1.
In Fig. 4, we show the distribution of $\mathrm{SNR}\geq 5$ hotspot candidates across the sky. Outside the mask, these appear roughly isotropically distributed, with an array of different $\eta_{\star}$ values. This is good: we do not find that all detection candidates cluster in regions of higher Galactic contamination. A more useful diagnostic is the distribution of SNR values, which is shown in Fig. 5. In total, we find $144$ candidates, with SNR up to $8$; it is important to note that our analysis has four free parameters ($\eta_{\star}$, $\eta_{\rm HS}$, and the hotspot center coordinates), thus moderately high SNR values are expected due to chance fluctuations. Furthermore, analysis of the sevem (commander) component-separated map finds only 48 (35) candidates, with a maximum detection significance of $6\sigma$. This indicates that the $\mathrm{SNR}\geq 6$ candidates (which all have $\mathrm{SNR}<5$ in component-separated maps) may be spurious instead of primordial.
@@
Tab. 1 lists the parameters of the six-frequency Planck $\mathrm{SNR}\geq 6$ hotspot candidates, with visualizations of the Planck maps around each point shown in Fig. 7. Each corresponds to a hotspot with temperature perturbation around $30\,\,\mu$K. Notably, all potential hotspots are small, with $\eta_{\star}<30\,\,\mathrm{Mpc}$ in all cases; as shown previously, physical hotspots of this size are the hardest to detect due to the increased noise on small-scales, and the broad Planck beam, which leads to small hotspots having the same spatial signatures as point sources – this is an important limitation of the method, which can be improved with the use of higher-resolution CMB data. Furthermore, six of the ten sources are located in a ring of radius $\sim 1^{\circ}$ and have similar inferred parameters, which suggests a ringing signature around a correlated source. Indeed, cross-matching with the public Planck catalogs (including point sources, Galactic cold clumps, tSZ clusters, and high-redshift sources), we find a non-thermal source (i.e., synchrotron) [50] at the center of the ring, which is clearly seen in Fig. 7. We additionally find a non-thermal source within $10^{\prime}$ of the $\mathrm{SNR}=6.0$ candidate and a $5\sigma$ point source within $13^{\prime}$ of the $\mathrm{SNR}=6.9$ candidate. All remaining sources are excluded by the Planck GAL020 mask (though this represents a harsh foreground cut), with the $\mathrm{SNR}=6.2$ candidate masked in the Planck inpainting map. All-in-all, we find no significant evidence for a hotspot with $\mathrm{SNR}\geq 6$ in the Planck data.
Figure 5: PDF of Planck PR4 detections at $\mathrm{SNR}\geq 5$. We compare results from the full six HFI channels (blue) and those from component-separated maps (orange, green) using two pipelines. These detections appear compatible with noise, given that the $\mathrm{SNR}\geq 6$ detections are not found in the component-separated pipeline.
@@ VI Summary & Discussion
Searching for localized features in the CMB is a task well-known to cluster cosmologists. Taking advantage of this parallel, we have utilized codes developed for thermal Sunyaev-Zel’dovich analysis to search for particle-production signatures in Planck PR4 data using matched-filter methods. It is difficult to overemphasize the simplicity of this approach: one simply replaces a tSZ cluster profile (and tSZ frequency dependence) with the desired primordial template (and blackbody frequency dependence). Extensive verification has demonstrated that this approach works: we can recover the input parameters of inflationary hotspots (both singlets and pairs) injected in realistic simulations, given sufficiently large couplings. From the Planck data, we find no evidence for new physics, with no robust detection of hotspots with $\mathrm{SNR}\geq 6$ (using a high threshold due to the four free parameters in the model). This yields a bound on shift-symmetry-violating couplings between the inflaton and a much heavier field, with $g\lesssim 10$ at time $\eta_{\star}>100\,\,\mathrm{Mpc}$. Assuming $60$ $e$-foldings of inflation between the horizon exit of a mode with physical size $1/H_{0}$ (the present day inverse Hubble parameter) and the end of inflation, $\eta_{\star}=100$ Mpc corresponds to $\approx 56$ $e$-foldings before the end of inflation.
@@
An important feature of the profile-finding analyses of this work is that their sensitivity does not depend on the number of hotspots present; we can detect any hotspot given sufficiently large $g,\eta_{\star}$. This differs significantly from power-spectrum-based analyses, whereupon the modification to $C_{\ell}^{TT}$ scales as $g^{2}N_{\rm spots}$ [e.g., 18]. In the right panel of Fig. 6, we compare the $5\sigma$ bound on $g$ from the two approaches, considering both one and one hundred hotspots. In the former case, the constraints from the localized profile-finding analysis dwarf those from the power spectrum, whilst in the latter case, searching for modifications to $C_{\ell}^{TT}$ may yield improved bounds, particularly at high $\eta_{\star}$. This indicates an important and generic point: rare events are best searched for via local-in-space analyses, whilst common events (giving some spectrum of sources) can be probed best with low-order correlators.
The analysis developed herein can be extended in a number of ways. Firstly, we have considered only temperature anisotropies; the Planck CMB dataset also contains polarization information, which can also be used to probe primordial hotspots since our mechanism produces adiabatic fluctuations that contribute to both $T$ and $E$. This will be particularly relevant for future surveys since large-scale temperature anisotropies are already cosmic-variance-limited. Furthermore, high-resolution data from the Atacama Cosmology Telescope [52, 53], South Pole Telescope [54, 55], Simons Observatory [56], and CMB-S4 [57] will facilitate analysis of smaller hotspots, whose spatial signatures are indistinguishable from point sources at the Planck resolution. This is shown in the right panel of Fig. 6, which gives the upper bounds on $g$ from an idealized experiment with neither a beam nor experimental noise for $\ell\in[2,3500]$. As shown, one could improve the limit on $g$ for small $\eta_{\star}$ by at least an order of magnitude in the future, though cosmic variance limits improvements at large $\eta_{\star}$. Finally, we note that our pipeline can be simply extended to a wide variety of other primordial features, for example those of [15, 20], allowing new insights into physics at the highest energies.
@@ Acknowledgements.
We thank Taegyun Kim and Neal Weiner for insightful comments and discussions, as well as the anonymous referee for insightful feedback. OHEP is a Junior Fellow of the Simons Society of Fellows and thanks Bandai Namco Entertainment Inc. for inspiring the NYC particle astrophysics and cosmology group meeting, as well as Netflix for motivating the title. SK was supported in part by NSF grant PHY-2210498 and the Simons Foundation. JCH acknowledges support from NSF grant AST-2108536, DOE grant DE-SC0011941, NASA grants 21-ATP21-0129 and 22-ADAP22-0145, the Sloan Foundation, and the Simons Foundation. This work utilized numpy [58], matplotlib [59], healpy [60], and HEALPix [61].
@@ Appendix A Visual Inspection of Planck Hotspot Candidates
In Fig. 7, we plot Cartesian projections of the Planck HFI frequency maps centered on the $\mathrm{SNR}\geq 6$ hotspot candidates given in Tab. 1. We also show the SNR map (given by $\hat{g}/\sigma(g)$, as defined in Eq. (6) with optimal $\eta_{\star},\eta_{\rm HS}$), computed for the $14.8^{\circ}\times 14.8^{\circ}$ tile in which the candidate was found. The first column appears consistent with a clump of dust (seen clearly at high frequency). Furthermore, we see a ringing feature for the third, which circles a synchrotron source identified in [50]. For the other sources, visual inspection does not yield definitive results.
Figure 7: SNR maps (top row) and frequency maps (other rows, with frequency indicated in GHz) displaying the $\mathrm{SNR}\geq 6$ hotspot candidates, which are circled in green. Each image is a $5^{\circ}\times 5^{\circ}$ Cartesian projection around the centers given in the title. The SNR maps have colorbar truncated to ±$\pm 5$, whilst the maximal value for each other plot is shown in purple. The bounding rectangles demarcate the flat-sky tiles used in the analysis, and the holes are due to point source masking.